OpenStudy (anonymous):

3x+2y=6 5x+3y=15 ~solve using Graphing ~Solve using Substitution ~Solve using Elimination ~Write an application problem

OpenStudy (lochana):

which part is hardest for you?

OpenStudy (anonymous):

graphing & elimination ( i don't know how to get a negative number instead of a postive with the numbers given) ~nothing times 3 gives -5~

OpenStudy (lochana):

okay. to graph equations, first you need to write equations for y \(y = \frac{6-3x}{2}\) is first one.

OpenStudy (lochana):

can you tell me second one? like I did for first one

OpenStudy (lochana):

\(y= \frac{15-5x}{3}\) yea?

OpenStudy (anonymous):

\[y=\frac{ 15-5 }{ 3 }\] .??

OpenStudy (lochana):

5x. not just 5

OpenStudy (anonymous):

why does it have to subbtract.?

OpenStudy (anonymous):

subtract*

OpenStudy (lochana):

\[3x +2y = 6 \\ 3x + 2y - 3x = 6 -3x \\ 2y = 6 -3x \\ \frac{2y}{2} = \frac{6-3x}{2} \\ y = \frac{6-3x}{2}\]

OpenStudy (lochana):

do same with 2nd one. you will get my answer

OpenStudy (anonymous):

to be honest i have no clue what you just did.. our teacher showed us an "easy" method.. which was... what times 3 gives us 6.. well 2, 3x2=6, so 2 was our x.. now what times 2 gives us 6.. well 3, 2x3=6, so 3 was our y..

OpenStudy (anonymous):

so our points were (2,3) & she graphed..

OpenStudy (lochana):

well, at the end, you have to graph them anyways. so these are the graphs 

OpenStudy (lochana):

so, solutions are points where 2 graphs meet. that's (0,3)

OpenStudy (anonymous):

well thanks for your help and time I appreciate it.. :)

OpenStudy (lochana):

you are welcome:)

OpenStudy (denisaboichuk):

c